# TextDiffuser-RL: Efficient and Robust Text Layout Optimization for High-Fidelity Text-to-Image Synthesis

### v1

@@ 1 Introduction
Figure 1: Overview of the Model Architecture, the diffusion-based synthesis process with GlyphEnv optimizing non-overlapping bounding box generation, layout planning and quality assessment based on OCR accuracy and spatial alignment.
@@ 2 Related Work
Reinforcement Learning with Human Feedback for Generative Models [16] explores reinforcement learning to fine-tune generative models based on human feedback. By using the classifier guidance, this approach enhances on the text-image alignment means, which allow precise control of the visual components regarding the textual annotation.
@@ 3.6 Training a Reinforcement Learning Algorithm with a Custom Environment
We trained three RL algorithms—Proximal Policy Optimization (PPO)[23] Deep Deterministic Policy Gradient (DDPG)[22], and Soft Actor-Critic (SAC)[29] within the GlyphEnv environment. All models operate in continuous action space, which is essential for our task as it allows fine-grained control over bounding box coordinates. This precision is necessary to minimize overlaps and adjust box positions smoothly. Using continuous actions lets the agent make small, precise changes rather than jumping between fixed steps. This helps the model learn better layouts that fit well within the canvas.
@@ 3.7 TextDiffuser-RL Model for Image Generation
In our TextDiffuser-RL model, we replace the original bounding box generation module with bounding boxes predicted by reinforcement learning (RL) agents, which provide optimized layouts that minimize overlap and improve text placement. The model pipeline begins with the tokenizer that encodes the input prompt into tokens. These tokens are then passed to the text_encoder to generate embeddings used by the diffusion process. The RL-generated bounding boxes serve as spatial constraints that guide the ddpm-scheduler[17], controlling the denoising steps during image generation. The U-Net model iteratively refines the latent representation by predicting and removing noise, while the vae decoder transforms the latent tensor into the final image. This integration of RL-based layouts into the model pipeline enables faster and more precise control over text placement in generated images.
@@ 4 Results
In this section, we present the evaluation of our proposed two-stage pipeline, which combines reinforcement learning (RL) for bounding box optimization with a diffusion-based image generation model. We describe the evaluation methodology, compare our method against state-of-the-art baselines, and showcase qualitative results. Evaluations are conducted on the MARIOEval benchmark and its LAIONEVal4000 subset.
@@ 4.1 Evaluation and Comparison
In this section, we present and analyze the results of our proposed two-stage pipeline, which integrates reinforcement learning for bounding box optimization with a diffusion-based image generation model. We evaluate the performance of different RL algorithms, assess the visual quality of generated images, and compare the inference efficiency of our approach.
@@ 4.1.1 Comparative Analysis of Reinforcement Learning Algorithms
To assess how well each reinforcement learning algorithm performs in optimizing text layout, we evaluated PPO, SAC, and DDPG on three key metrics: mean reward, standard deviation of rewards, and average IoU overlap between bounding boxes. Among the three, PPO showed the most promising results with a mean reward of 579.54 and an average IoU overlap of 0.26, meaning it consistently generated layouts with minimal overlap. In contrast, SAC and DDPG performed noticeably worse. SAC produced a much lower mean reward of -4222.90 and a higher IoU overlap of 0.61, indicating more frequent and larger overlaps between boxes. DDPG performed the worst overall, with a mean reward of -6200.37 and the highest IoU overlap of 0.70.
@@
These results suggest that PPO is more effective at learning to place bounding boxes with minimal intersection, which is critical for downstream tasks like image generation. The high variance in rewards for SAC and DDPG also indicates instability during training.
@@ 4.1.2 Text-to-Image Generation Quality Analysis
To evaluate the visual and textual quality of the generated images, we used CLIPScore[21] and OCR metrics[32]. CLIPScore measures the semantic similarity between the input text prompt and the generated image by comparing their embeddings using the CLIP model. This metric helps assess how well the image reflects the intended textual content. On the other hand, OCR accuracy evaluates how accurately the generated text can be extracted from the image using an optical character recognition (OCR) engine. In addition to these metrics, we used the MARIOEval[27] benchmark, which consists of 5,414 image-description pairs designed for evaluating multimodal generation quality.
@@
SA-OcrPaint [28] 85.70 66.95 35.70
ARTIST [30] 86.79 73.73 34.82
TextDiffuser-RL (Ours) 71.61 32.56 34.73
@@
TextDiffuser and TextDiffuser-2 exhibit strong performance in OCR and CLIPScore, benefiting from LM-based layout generation. While our RL-based approach does not rely on such heavy architectures, it still achieves competitive OCR F1 (71.61) and CLIPScore (34.73), demonstrating the effectiveness of reinforcement learning for layout optimization. Compared to SA-OcrPaint, which attains the highest CLIPScore (35.70) and strong OCR metrics, our model achieves nearly comparable accuracy with drastically lower resource demands. This makes TextDiffuser-RL an efficient alternative without significant compromise in performance.
@@ 5 Discussion
Our experimental results demonstrate that reinforcement learning, specifically PPO, significantly improves bounding box layout generation in terms of overlap reduction and inference efficiency. This has a strong positive impact on the downstream diffusion-based image generation, as seen in the improved qualitative results and competitive CLIPScore. One possible reason PPO outperforms SAC and DDPG is its clipped surrogate objective, which stabilizes training in our sparse reward environment. In contrast, the off-policy nature of SAC and DDPG may lead to unstable updates when facing continuous spatial tasks with strong constraints.
While our method performs well in general, it struggles with extremely dense or long text prompts, occasionally producing overlapping boxes. These edge cases may benefit from hybrid approaches combining RL with heuristics or learned priors. Additionally, the current system does not follow a strict sentence order during generation and is currently limited to rectangular shapes.
Future work could focus on integrating character-aware models into the RL reward function, expanding to multilingual datasets, or incorporating user-editable layout guidance. We also aim to extend the system to support non-rectangular shape generation, which could enhance versatility and impact in the graphic design industry.

### v2

@@ 1 Introduction
Figure 1: Overview of the Model Architecture, the diffusion-based synthesis process with GlyphEnv optimizing non-overlapping bounding box generation, layout planning, and quality assessment based on OCR accuracy and spatial alignment.
@@ 2 Related Work
Reinforcement Learning with Human Feedback for Generative Models [16] explores reinforcement learning to fine-tune generative models based on human feedback. By using the classifier guidance, this approach builds on the text-image alignment means, which allow precise control of the visual components regarding the textual annotation.
@@ 3.6 Training a Reinforcement Learning Algorithm with a Custom Environment
We trained three RL algorithms—Proximal Policy Optimization (PPO)[23], Deep Deterministic Policy Gradient (DDPG)[22], and Soft Actor-Critic (SAC)[29] within the GlyphEnv environment. All models operate in continuous action space, which is essential for our task as it allows fine-grained control over bounding box coordinates. This precision is necessary to minimize overlaps and adjust the box positions smoothly. Using continuous actions lets the agent make small, precise changes rather than jumping between fixed steps. This helps the model learn better layouts that fit well within the canvas.
@@ 3.7 TextDiffuser-RL Model for Image Generation
In our TextDiffuser-RL model, we replace the original bounding box generation module with bounding boxes predicted by reinforcement learning (RL) agents, which provide optimized layouts that minimize overlap and improve text placement. The model pipeline begins with the tokenizer that encodes the input prompt into tokens. These tokens are then passed to the text_encoder to generate embeddings used by the diffusion process. RL-generated bounding boxes serve as spatial constraints that guide the ddpm_scheduler [17], controlling the denoising steps during image generation. The U-Net model iteratively refines the latent representation by predicting and removing noise, while the vae decoder transforms the latent tensor into the final image. This integration of RL-based layouts into the model pipeline enables faster and more precise control over text placement in generated images.
@@ 4 Results
In this section, we present the evaluation of our proposed two-stage pipeline, which combines reinforcement learning (RL) for bounding-box optimization with a diffusion-based image generation model. We describe the evaluation methodology, compare our method with state-of-the-art baselines, and present qualitative results. Evaluations are performed on the MARIOEval benchmark and its LAIONEVal4000 subset.
@@ 4.1 Evaluation and Comparison
In this section, we present and analyze the results of our proposed two-stage pipeline, which integrates reinforcement learning for bounding box optimization with a diffusion-based image generation model. We evaluated the performance of different RL algorithms, evaluated the visual quality of generated images, and compared the inference efficiency of our approach.
@@ 4.1.1 Comparative Analysis of Reinforcement Learning Algorithms
To assess how well each reinforcement learning algorithm performs in optimizing text layout, we evaluated PPO, SAC, and DDPG on three key metrics: mean reward, standard deviation of rewards, and average IoU overlap between bounding boxes. Among the three, PPO showed the most promising results with a mean reward of 579.54 and an average IoU overlap of 0.26, which means it consistently generated layouts with minimal overlap. In contrast, SAC and DDPG performed noticeably worse. SAC produced a much lower mean reward of -4222.90 and a higher IoU overlap of 0.61, indicating more frequent and larger overlaps between boxes. DDPG performed the worst overall, with a mean reward of -6200.37 and the highest IoU overlap of 0.70.
@@
These results suggest that PPO is more effective in learning to place bounding boxes with minimal intersection, which is critical for downstream tasks such as image generation. The high variance in rewards for SAC and DDPG also indicates instability during training.
@@ 4.1.2 Text-to-Image Generation Quality Analysis
To evaluate the visual and textual quality of the generated images, we used CLIPScore[21] and OCR metrics[32]. CLIPScore measures the semantic similarity between the input text prompt and the generated image by comparing their embeddings using the CLIP model. This metric helps assess how well the image reflects the intended textual content. However, OCR accuracy evaluates how accurately the generated text can be extracted from the image using an optical character recognition (OCR) engine. In addition to these metrics, we used the MARIOEval[27] benchmark, which consists of 5,414 image-description pairs designed to evaluate the quality of multimodal generation.
@@
TextDiffuser-RL (Ours) 69.91 34.30 34.47
@@
TextDiffuser and TextDiffuser-2 exhibit strong performance in OCR and CLIPScore, benefiting from LM-based layout generation. While our RL-based approach does not rely on such heavy architectures, it still achieves competitive OCR F1 (69.91) and CLIPScore (34.47), demonstrating the effectiveness of reinforcement learning for layout optimization. Compared to SA-OcrPaint, which attains the highest CLIPScore (35.70) and strong OCR metrics, our model achieves nearly comparable accuracy with drastically lower resource demands. This makes TextDiffuser-RL an efficient alternative without a significant performance compromise.
@@ 5 Discussion
Our experimental results demonstrate that reinforcement learning—particularly PPO—offers a compelling approach for bounding box layout generation, significantly reducing overlap while maintaining high inference efficiency. This improvement positively influences downstream diffusion-based image synthesis, as reflected by our competitive CLIPScore and qualitative outputs. A likely reason PPO outperforms SAC and DDPG lies in its clipped surrogate objective, which provides greater stability during training in environments with sparse and delayed rewards. In contrast, the off-policy nature of SAC and DDPG appears less suited to our continuous, spatially constrained task.
Despite these strengths, our method has several limitations. First, the current environment is static: it generates a fixed number of bounding boxes and assumes uniform box sizes, regardless of prompt length or spatial context. This restricts flexibility when dealing with varying prompt complexities. Additionally, in edge cases involving dense or unusually long text prompts, the model may still produce overlapping boxes. The system also does not enforce a strict sentence or semantic ordering, which can limit coherence in generated layouts.
Future work may explore dynamic environments that allow variable numbers and sizes of boxes, possibly conditioned on the prompt’s linguistic structure. Combining PPO with heuristic refinements or learned priors may further improve robustness. Other promising directions include incorporating character-aware reward signals, supporting multilingual prompts, enabling non-rectangular shape layouts, and adding user-controllable layout editing tools to make the system more practical for real-world design tasks.